A manufacturer ships toasters in cartons of 20. In each​ carton, they estimate a 10​% chance that any one of the toasters will need to be sent back for minor repairs. What is the probability that in a​ carton, there will be exactly 2 toasters that need​ repair?

The probability of exactly 2 toasters needing repair is _____
Answer #1

solution:

Given that ,

p = 10%=0.10

1 - p = 1 - 0.10 = 0.9

n =20

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X =2 ) = (20 C 2) * 0.102 * (0.9) 20- 2

probability =0.2852